# Why Study the Spherical Convexity of Non-Homogeneous Quadratic Functions, and What Makes It Surprising?

### v1

@@ Abstract
This paper presents necessary, sufficient, and equivalent conditions for the spherical convexity of non-homogeneous quadratic functions. In addition to motivating this study and identifying useful criteria for determining whether such functions are spherically convex, we discovered surprising properties that distinguish spherically convex quadratic functions from their geodesically convex counterparts in both hyperbolic and Euclidean spaces. Since spherically convex functions over the entire sphere are constant, we restricted our focus to proper spherically convex subsets of the sphere. Although most of our results pertain to non-homogeneous quadratic functions on the spherically convex set of unit vectors with positive coordinates, we also present findings for more general spherically convex sets. Beyond the general non-homogeneous quadratic functions, we consider explicit special cases where the matrix in the function’s definition is of a specific type, such as positive, diagonal, and Z-matrix.
@@ 2 Basic terminology
A set $K\subseteq\mathbb{R}^{n}$ is called cone if $\lambda x\in K$, for all $x\in K$ and all $\lambda>0$. A cone $K\subseteq\mathbb{R}^{n}$ is called convex cone if is a if $x+y\in K$ for any $x,y\in K$. It follows that a cone $K\subseteq\mathbb{R}^{n}$ is a convex cone if and only if it is a convex set. A cone $K\subseteq\mathbb{R}^{n}$ is called pointed cone if there is no $x\in K\setminus\{0\}$ such that $-x\in K$. A cone $K\subseteq\mathbb{R}^{n}$ is called closed cone if it is a closed set. A cone $K\subseteq\mathbb{R}^{n}$ is called proper if it is a pointed closed convex cone with nonempty interior. The polar $K^{\perp}$ of a cone $K\subseteq\mathbb{R}^{n}$ is defined by $K^{\perp}=\{x\in\mathbb{R}^{n}:\left\langle x,y\right\rangle\geq 0,\,\forall y\in K\}$.
@@ 4 Main results
Throughout this paper we restrict ourselves to proper cones $K\subseteq\mathbb{R}^{n}$, that is to pointed closed convex cones with nonempty interior. Denote the dual of $K$ by $K^{*}$ and the polar of $K$ by $K^{\perp}$. Let $c\in\mathbb{R}$, $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix and $b\in\mathbb{R}^{n}$. From now on $f$ will always denote a function $f:\mathbb{S}^{n}\cap\operatorname{int}(K)\to\mathbb{R}$ defined by
@@
(ii)$\implies$(iv): Let any $u,v\in\mathbb{S}^{n}$ such that $v\in K$. Denote by $v^{2}\in K^{\oplus}$ the vector of coordinates $(v^{i})^{2}$ and by $u^{2}\in K^{\oplus}$ the vector of coordinates $(u_{i})^{2}$. Since $v\in\mathbb{S}^{n}\cap K$, we have $v-v^{2}\in K=K^{\oplus}$. Thus, the copositivity of $C$, $-b\in K=K^{\oplus}$ and $v-v^{2}\in K^{\oplus}$ imply

### v2

# Why Study Spherical Convexity of Non-Homogeneous Quadratic Functions? What Makes It Surprising?

@@ Abstract
This paper establishes necessary, sufficient, and equivalent conditions for the spherical convexity of non-homogeneous quadratic functions. By examining criteria for determining spherical convexity, we identified unique properties that differentiate spherically convex quadratic functions from their geodesically convex counterparts in both hyperbolic and Euclidean spaces. Since spherically convex functions over the entire sphere are constant, our analysis focuses on proper spherically convex subsets of the sphere. Our primary results concern non-homogeneous quadratic functions on the spherically convex set of unit vectors with positive coordinates. We also extend our findings to more general spherically convex sets. Additionally, the paper explores special cases of non-homogeneous quadratic functions where the defining matrix is of a specific type, such as positive, diagonal, or a Z-matrix. This study not only provides useful criteria for spherical convexity but also reveals surprising characteristics of spherically convex quadratic functions, contributing to a deeper understanding of convexity in spherical geometries.
@@ 2 Basic terminology
A set $K\subseteq\mathbb{R}^{n}$ is called cone if $\lambda x\in K$, for all $x\in K$ and all $\lambda>0$. A cone $K\subseteq\mathbb{R}^{n}$ is called convex cone if is a if $x+y\in K$ for any $x,y\in K$. It follows that a cone $K\subseteq\mathbb{R}^{n}$ is a convex cone if and only if it is a convex set. A cone $K\subseteq\mathbb{R}^{n}$ is called pointed cone if there is no $x\in K\setminus\{0\}$ such that $-x\in K$. A cone $K\subseteq\mathbb{R}^{n}$ is called closed cone if it is a closed set. A cone $K\subseteq\mathbb{R}^{n}$ is called proper if it is a pointed closed convex cone with nonempty interior. The dual $K^{\perp}$ of a cone $K\subseteq\mathbb{R}^{n}$ is defined by
$K^{*}=\{x\in\mathbb{R}^{n}:\left\langle x,y\right\rangle\geq 0,\,\forall y\in K\}.$
The polar $K^{\perp}$ of $K$ is defined by
$K^{\perp}=\{x\in\mathbb{R}^{n}:\left\langle x,y\right\rangle\leq 0,\,\forall y\in K\}.$
@@ 4 Main results
Throughout this paper we restrict ourselves to proper cones $K\subseteq\mathbb{R}^{n}$, that is to pointed closed convex cones with nonempty interior. Let $c\in\mathbb{R}$, $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix and $b\in\mathbb{R}^{n}$. From now on $f$ will always denote a function $f:\mathbb{S}^{n}\cap\operatorname{int}(K)\to\mathbb{R}$ defined by
@@
(ii)$\implies$(iv): Let any $u,v\in\mathbb{S}^{n}$ such that $v\in K$. Denote by $v^{2}\in K^{\*}$ the vector of coordinates $(v^{i})^{2}$ and by $u^{2}\in K^{\*}$ the vector of coordinates $(u_{i})^{2}$. Since $v\in\mathbb{S}^{n}\cap K$, we have $v-v^{2}\in K=K^{\*}$. Thus, the copositivity of $C$, $-b\in K=K^{\*}$ and $v-v^{2}\in K^{\*}$ imply